# Can We Verify Step by Step for Incorrect Answer Detection?

### v2

@@ A.1 Dataset Processing
  • GSM8K (Cobbe et al., 2021) contains a collection of grade school math word problems. Our research uses the entire test set. MIT license: GSM8K license.
  • MATH (Hendrycks et al., 2021) consists of competition-level mathematics problems. For our study, we conducted a random sampling of 1,000 problems from its test set, which contains a total of 5,000 examples. MIT license: MATH license.
  • StrategyQA (Geva et al., 2021) is a question answering benchmark that focuses on open-domain questions where the required reasoning steps are implicit in the question and should be inferred using a strategy. We use the open-domain setting (question-only set) from BIG-bench Collaboration (Srivastava et al., 2022), and randomly select 1000 examples from its test set. MIT license: StrategyQA license.
  • Play originates from BIG-bench (Srivastava et al., 2022), designed to assess the model’s ability to deduce whether two closely sequenced "lines" from a dialogue in a Shakespearean play were spoken by the same character or by different characters. This particular task challenges the model’s understanding of context, character consistency, and dialogue progression within the intricate settings of Shakespearean works. We used the entire test set. Link: Play link.
  • Physics (from BIG-bench (Srivastava et al., 2022)), is crafted to evaluate a model’s understanding of physics by requiring it to identify the appropriate formula for solving physics word problems. This dataset comprises 229 multiple-choice questions. For our study, we used the first two questions as few-shot demonstrations, and the remaining questions serve as the test set. Link: Physics link.
  • The fact verification dataset, FEVER (Thorne et al., 2018), categorizes claims as "SUPPORTS," "REFUTES," or "NOT ENOUGH INFO" by analyzing evidence paragraphs obtained from Wikipedia. The data annotations of FEVER utilize information from Wikipedia, which is obtained in accordance with the Wikipedia Copyright Policy. We use the FEVER subset by Zhao et al. (2023a), where 1,000 of 3,000 samples are randomly selected to balance the number of right and wrong predictions in CoT.
  • Adversarial HotpotQA (Yang et al., 2018) is a multi-hop question answering dataset. We adopt the subset proposed by Ye and Durrett (2022), which presents a challenging set of data in which the correct and incorrect predictions are evenly balanced through the use of their model.
  • 2WikiMultihop (Ho et al., 2020) is a dataset designed for multi-hop question-answering tasks. It leverages the structured format found in Wikidata and utilizes logical rules to provide answers. It is released under the Apache License 2.0, allowing users to modify and use it for research purposes. We use the subset provided by Zhao et al. (2023a), which is a subsample with size 1000 out of the dev set of 12,576 samples.
@@
We exclude open-ended generation tasks, such as text summarization, from our R2PE benchmark. These tasks can use proxy metrics (Lin, 2004; Ganesan, 2018; Yuan et al., 2021) to measure performance; however, it is difficult to label responses as true or false directly. A potential solution is to calculate the Pearson correlation between metrics and scores that capture the quality of reasoning chains, and we leave this to future work.
@@ A.2 Queried LLMs
  • GPT-3 (Brown et al., 2020), where we specifically use the text-davinci-003 engine. This engine became deprecated on 1 January 2024.
@@
When examining performance in open-domain question-answering tasks, we replaced GPT-3.5-turbo with GPT-4 (OpenAI, 2023). The reason for this replacement was the observed reluctance of GPT-3.5-turbo to engage with questions from these datasets, often defaulting to responses indicating a lack of sufficient information. Attempts to solicit responses from the GPT-3.5-turbo model frequently resulted in non-committal replies such as ’I do not have enough information,’ also documented by Zheng et al. (2023) and illustrated through examples in Appendix E. We hypothesize that this behavior may be attributable to the removal of contexts, a characteristic of the open-domain nature of these tasks used in our study.
@@ A.3 Extraction and Manual Cleaning
In the MATH dataset, the answers are clearly demarcated with \$\textbackslash boxed\{\}$, which facilitates straightforward automated extraction. Specifically, we use the Python codes provided by (Zheng et al., 2023) to parse the answers. Two questions within (MATH from text-davici-003) and one quesion in (MATH from Mixtral-8x7b) are discarded due to parsing errors induced by the aforementioned code. In addition, a thorough manual inspection is conducted to address and correct the formatting in some unconventional cases. This was particularly necessary for certain questions where the responses deviated from standard sentential structures. The number of such cases is given in Table 8. Some of these examples are provided in Table 9.
@@
Q: If you meet a thief, you may suspect him, by virtue of your office, to be no true man; and, for such kind of men, the less you meddle or make with them, why the more is for your honesty. If we know him to be a thief, shall we not lay hands on him? Truly, by your office, you may; but I think they that touch pitch will be defiled: the most peaceable way for you, if you do take a thief, is to let him show himself what he is and steal out of your company. In the preceding dialogue, were the lines "Truly, by your office, you may; but I think they that touch pitch will be defiled: the most peaceable way for you, if you do take a thief, is to let him show himself what he is and steal out of your company." and "If we know him to be a thief, shall we not lay hands on him?" spoken by the same person or different people?
@@
In a manner akin to the StrategyQA dataset, each response within the Play dataset is anticipated to be classified as either ’different’ or ’same’. To ensure uniformity, we employed an identical pipeline for the extraction of these answers. And 16 questions are filtered in the (Play from Gemini Pro). The identified answer trigger words for the Play dataset and the number of examples that are manually corrected are listed in Table 7 and  8 correspondingly.
@@
For the FEVER dataset, which includes a neutral class labeled "NOT ENOUGH INFO," we optimized our processing pipeline by assigning this label to responses deviating from the designated prompt format. Unlike FEVER, both the HotpotQA and 2WikiMultihop datasets comprise free-form answers, avoiding the need for a manual cleaning phase. We used Python scripts from Zhao et al. (2023a) to extract answers and rationales.
@@ Appendix B More Examples of R2PE
Question $Q$ The question or claim that is used to query LLM from the original dataset. What was the nationality of the actor playing a Russian hostage taker in "Hostage for a Day"?
@@
Response The reponses generated by "LLM Name" to answer Question $Q$, containing five responses with reasoning chains and answers $R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})$ First, “Hostage for a Day" is a 1994 TV movie directed by John Candy. Second, the Russian hostage-taker character, Boris, is portrayed by George Wendt, an American actor. The answer is American. …
@@ C.2 Precision and Recall
In Table 1, it is evident that PDS consistently surpasses ADS regarding F1 score. To further elucidate these findings, we also present precision and recall as supplementary metrics in Table 15 and 16. While the ADS exhibits higher precision in most cases, the PDS significantly enhances recall across all subgroups. On average, PDS reduces precision by $5.72\%$ and increases recall by $15.68\%$, suggesting a better balance between precision and recall, culminating in an enhanced F1 score. From practical respect, where the detection of all incorrect instances is paramount, even at the cost of occasionally misidentifying correct answers as potentially incorrect, a higher recall is preferable due to its emphasis on comprehensive falsehood detection.
@@ C.3 About the Threshold
Following Li et al. (2023), we raise the threshold $H$ for discriminative tasks where the answers are from a fixed set and often have greater answer agreement among responses. Such tasks include binary (yes/no) choices or multiple-choice queries. More concretely, the threshold $H$ for ADS is set to be 4.5 and $0.4$ for PDS for subsets in R2PE associated with StrategyQA, Play, Physics, and FEVER. Despite this change, our thresholds are still not task-specific. That is because we can only in advance whether a query is a discriminative one. In our previous example, if the user asks a yes/no or multiple choice question, we set the threshold to $4.5$ for ADS and $0.4$ for PDS, respectively. Otherwise, we set the threshold to $2.5$ for ADS and 00 for PDS.
@@
We also show the effect of varying the threshold of PDS in Figure 5.
@@
C.4 HaloCheck and SelfCheckNLI
HaloCheck (Elaraby et al., 2023) and SelfCheckNLI (Manakul et al., 2023) are two hallucination detection methods for open-ended generation, which can serve as alternative implementations of PSS. HaloCheck uses the same formula of PPSS as PDS-avg, albeit calculating the average over only a subset of all response pairs: $PSS_{HaloCheck}=mean_{i<j}PPSS(R_{i},R_{j})$. It is important to note that $PPSS(A,B)$ is not symmetric with respect to A and B, introducing a nuanced difference in results. In contrast, SelfCheckNLI adopts a distinct PPSS implementation, leveraging DeBERTa-v3-large (He et al., 2021) finetuned to MNLI (Williams et al., 2018). Additionally, SelfCheckNLI also alters the PSS formulation $PSS_{SelfCheckNLI}=mean_{j}PPSS(R,R_{j})$, where $R$ represents a randomly selected response from the set of highly agreeable answers. This method diverges from the full utilization of all responses by focusing the evaluation on the factuality of a single generated response, as highlighted by Elaraby et al. (2023). This deviation from utilizing the collective information present in all responses leads to a notable performance decrease when replacing our PDS implementation with that of PDS-SelfCheckNLI, as evidenced by the results in Table 3.
@@ Appendix D Combining with verify-and-edit
VE (Zhao et al., 2023a) is a two-stage framework to edit unreliable reasoning chains. The first stage is to decide when to edit. The original VE uses the answer consistency to indicate when to edit. Specifically, if the majority cannot agree on the same answer, it will be revised. The second stage is about how to edit the selected rationales. This stage contains three steps: verifying question generation, context retrieval, and answering again.
To edit the rationale $r=s_{1},...,s_{n}$, we first generate one question $q_{i}$ for each sentence $s_{i}$ via prompting. The question $q_{i}$ is to verify the sentence $s_{i}$ in the rationale $r$ and is used as a query to retrieve the relevant context $c_{i}$. $c_{i}$ is composed of the top 3 sentences most similar to query $q_{i}$ ranked by the pre-trained Sentence BERT model (Reimers and Gurevych, 2019) in the retrieved documents. Based on the retrieved context $c_{i}$, the answer for verifying question $q_{i}$ is generated, denoted by $s^{\prime}_{i}$. Finally, we have a new rationale $r^{\prime}=s^{\prime}_{1},...,s^{\prime}_{n}$ that is edited leveraging external knowledge. The prompts for verifying question generation and answering again are provided in Table 17,  18, and  19. Specifically, we use the orical setting from Zhao et al. (2023a) to counteract the influence of retriever performance.
@@
Alberta Ferretti (Cattolica, 1950) is an Italian fashion designer and dressmaker. Her showroom is in Milan, Italy but her studio is in the village of Cattolica, near Rimini, Italy. Rimini (] ; Romagnol dialect: "Rémin"; Latin: "Ariminum") is a city of 146,606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Rimini. Queequeg (] ; Romagnol dialect: "Rémin"; Latin: "Ariminum") is a city of 546606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Queequeg. Chinatown (] ; Romagnol dialect: "Rémin"; Latin: "Ariminum") is a city of 346606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Chinatown .
@@
Appendix E Examples of GPT-3.5-turbo and mistral-medium
During the querying phase, we try to use GPT-3.5-turbo to answer questions from HotpotQA and 2WikiMultihop. However, it always outputs responses like "I’m sorry, I do not have enough information…". We assume this recurring outcome is attributed to the open-domain nature of these tasks. Examples are given in Table 20. Therefore, we use GPT-4 as a substitute.

### v4

@@ A.1 Dataset Processing
  • GSM8K [Cobbe et al., 2021] contains a collection of grade school math word problems. Our research uses the entire test set. MIT license: GSM8K license.
  • MATH [Hendrycks et al., 2021] consists of competition-level mathematics problems. For our study, we conducted a random sampling of 1,000 problems from its test set, which contains a total of 5,000 examples. MIT license: MATH license.
  • StrategyQA [Geva et al., 2021] is a question answering benchmark that focuses on open-domain questions where the required reasoning steps are implicit in the question and should be inferred using a strategy. We use the open-domain setting (question-only set) from BIG-bench Collaboration [Srivastava et al., 2022], and randomly select 1000 examples from its test set. MIT license: StrategyQA license.
  • Play originates from BIG-bench [Srivastava et al., 2022], designed to assess the model’s ability to deduce whether two closely sequenced ”lines” from a dialogue in a Shakespearean play were spoken by the same character or by different characters. This particular task challenges the model’s understanding of context, character consistency, and dialogue progression within the intricate settings of Shakespearean works. We used the entire test set. Link: Play link.
  • Physics (from BIG-bench [Srivastava et al., 2022]), is crafted to evaluate a model’s understanding of physics by requiring it to identify the appropriate formula for solving physics word problems. This dataset comprises 229 multiple-choice questions. For our study, we used the first two questions as few-shot demonstrations, and the remaining questions serve as the test set. Link: Physics link.
  • The fact verification dataset, FEVER (Thorne et al., 2018), categorizes claims as ”SUPPORTS,” ”REFUTES,” or ”NOT ENOUGH INFO” by analyzing evidence paragraphs obtained from Wikipedia. The data annotations of FEVER utilize information from Wikipedia, which is obtained in accordance with the Wikipedia Copyright Policy. We use the FEVER subset by Zhao et al. [2023], where 1,000 of 3,000 samples are randomly selected to balance the number of right and wrong predictions in CoT.
  • Adversarial HotpotQA [Yang et al., 2018] is a multi-hop question answering dataset. We adopt the subset proposed by Ye and Durrett [2022], which presents a challenging set of data in which the correct and incorrect predictions are evenly balanced through the use of their model.
  • 2WikiMultihop [Ho et al., 2020] is a dataset designed for multi-hop question-answering tasks. It leverages the structured format found in Wikidata and utilizes logical rules to provide answers. It is released under the Apache License 2.0, allowing users to modify and use it for research purposes. We use the subset provided by Zhao et al. [2023], which is a subsample with size 1000 out of the dev set of 12,576 samples.
@@
We exclude open-ended generation tasks, such as text summarization, from our R2PE benchmark. These tasks can use proxy metrics [Lin, 2004; Ganesan, 2018] to measure performance; however, it is difficult to label responses as true or false directly. A potential solution is to calculate the Pearson correlation between metrics and scores that capture the quality of reasoning chains, and we leave this to future work.
@@ A.2 Queried LLMs
  • GPT-3, where we specifically use the text-davinci-003 engine. This engine became deprecated on 1 January 2024.
@@
When examining performance in open-domain question-answering tasks, we replaced GPT-3.5-turbo with GPT-4 [OpenAI, 2023]. The reason for this replacement was the observed reluctance of GPT-3.5-turbo to engage with questions from these datasets, often defaulting to responses indicating a lack of sufficient information. Attempts to solicit responses from the GPT-3.5-turbo model frequently resulted in non-committal replies such as ’I do not have enough information,’ also documented by Zheng et al. [2023] and illustrated through examples in Appendix F. We hypothesize that this behavior may be attributable to the removal of contexts, a characteristic of the open-domain nature of these tasks used in our study.
@@ A.3 Extraction and Manual Cleaning
In the MATH dataset, the answers are clearly demarcated with \$\textbackslash boxed\{\}$, which facilitates straightforward automated extraction. Specifically, we use the Python codes provided by [Zheng et al., 2023] to parse the answers. Two questions within (MATH from text-davici-003) and one quesion in (MATH from Mixtral-8x7b) are discarded due to parsing errors induced by the aforementioned code. In addition, a thorough manual inspection is conducted to address and correct the formatting in some unconventional cases. This was particularly necessary for certain questions where the responses deviated from standard sentential structures. The number of such cases is given in Table 10. Some of these examples are provided in Table 11.
@@
Q: If you meet a thief, you may suspect him, by virtue of your office, to be no true man; and, for such kind of men, the less you meddle or make with them, why the more is for your honesty. If we know him to be a thief, shall we not lay hands on him? Truly, by your office, you may; but I think they that touch pitch will be defiled: the most peaceable way for you, if you do take a thief, is to let him show himself what he is and steal out of your company. In the preceding dialogue, were the lines ”Truly, by your office, you may; but I think they that touch pitch will be defiled: the most peaceable way for you, if you do take a thief, is to let him show himself what he is and steal out of your company.” and ”If we know him to be a thief, shall we not lay hands on him?” spoken by the same person or different people?
@@
In a manner akin to the StrategyQA dataset, each response within the Play dataset is anticipated to be classified as either ’different’ or ’same’. To ensure uniformity, we employed an identical pipeline for the extraction of these answers. And 16 questions are filtered in the (Play from Gemini Pro). The identified answer trigger words for the Play dataset and the number of examples that are manually corrected are listed in Table 9 and  10 correspondingly.
@@
For the FEVER dataset, which includes a neutral class labeled ”NOT ENOUGH INFO,” we optimized our processing pipeline by assigning this label to responses deviating from the designated prompt format. Unlike FEVER, both the HotpotQA and 2WikiMultihop datasets comprise free-form answers, avoiding the need for a manual cleaning phase. We used Python scripts from Zhao et al. [2023] to extract answers and rationales.
@@ Appendix B More Examples of R2PE
Question $Q$ The question or claim that is used to query LLM from the original dataset. What was the nationality of the actor playing a Russian hostage taker in ”Hostage for a Day”?
@@
Response The reponses generated by ”LLM Name” to answer Question $Q$, containing five responses with reasoning chains and answers $R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})$ First, “Hostage for a Day” is a 1994 TV movie directed by John Candy. Second, the Russian hostage-taker character, Boris, is portrayed by George Wendt, an American actor. The answer is American. …
@@ C.2 Precision and Recall
In Table 1, it is evident that PDS consistently surpasses ADS regarding F1 score. To further elucidate these findings, we also present precision and recall as supplementary metrics in Table 17 and 18. While the ADS exhibits higher precision in most cases, the PDS significantly enhances recall across all subgroups. On average, PDS reduces precision by $5.72\%$ and increases recall by $15.68\%$, suggesting a better balance between precision and recall, culminating in an enhanced F1 score. From practical respect, where the detection of all incorrect instances is paramount, even at the cost of occasionally misidentifying correct answers as potentially incorrect, a higher recall is preferable due to its emphasis on comprehensive falsehood detection.
@@ C.3 About the Threshold
Following Li et al. [2023], we raise the threshold $H$ for discriminative tasks where the answers are from a fixed set and often have greater answer agreement among responses. Such tasks include binary (yes/no) choices or multiple-choice queries. More concretely, the threshold $H$ for ADS is set to be 4.5 and $0.4$ for PDS for subsets in R2PE associated with StrategyQA, Play, Physics, and FEVER. Despite this change, our thresholds are still not task-specific. That is because we can only in advance whether a query is a discriminative one. In our previous example, if the user asks a yes/no or multiple choice question, we set the threshold to $4.5$ for ADS and $0.4$ for PDS, respectively. Otherwise, we set the threshold to $2.5$ for ADS and 00 for PDS.
@@
We also show the effect of varying the threshold of PDS in Figure 5. To adjust the threshold according to the number of choices in a question, we demonstrate the impact of varying the threshold of ADS on StrategyQA (2 options), FEVER (3 options), and Physics (4 options) in Table 19. The results indicate that it is suitable to set the threshold of ADS to 4.5 for discriminative tasks. The threshold for PDS is then adjusted to 0.4 for these tasks according to equation (1) ((4.5 - 2.5) / 2.5 = 0.4).
@@
dataset/threshold 2.5 3.5 4.5
StrategyQA 8.88 44.89 57.51
Fever 6.47 35.38 53.01
Physics 8.11 40.00 58.65
Table 19: Accuracy across different thresholdsl of ADS.
C.4 Detailed Setting of Ablation Study
To understand the contribution of PDS components to the overall performance, we have carried out an extensive ablation study involving several variants in Section 4.2. We here list the formulas for each of the variants.
PDS-ADS. we remove the $\overline{ADS}$ in our Equation (1), and the alternative of PDS becomes
[…]
$=\text{mean}_{i\neq j}\text{PPSS}(R_{i},R_{j})$
$=\text{mean}_{i\neq j}\min_{j}\max_{i}\bigl{(}ent(a_{i},b_{j})-con(a_{i},b_{j})\bigr{)}.$
PDS w/o ans. The original PSS considers all $PPSS(R_{i},R_{j})$, where each response $R_{i}$ equals a reasoning path $r_{i}$ and an answer $a_{i}$. As the information from the answers $a_{i}$’ is already used to obtain the final answer $a$, we can discard them and only compute the entailment score among rationale pairs $PPSS(r_{i},r_{j})$, that is,
$PDS_{w/o\>ans}=(\overline{ADS}+\text{mean}_{i\neq j}PPSS(r_{i},r_{j}))/2$
$=(\overline{ADS}+\text{mean}_{i\neq j}\min_{j}\max_{i}\bigl{(}ent(a_{i},b_{j})-con(a_{i},b_{j})\bigr{)})/2.$
$\text{PDS}=(\overline{ADS}+\text{mean}_{i\neq j}PPSS(r_{i},r_{j}))/2$
$=(\overline{ADS}+\text{mean}_{i\neq j}\text{avg}_{j}\text{max}_{i}(ent(a_{i},b_{j})-con(a_{i},b_{j})))/2.$
C.5 HaloCheck and SelfCheckNLI
HaloCheck [Elaraby et al., 2023] and SelfCheckNLI [Manakul et al., 2023] are two hallucination detection methods for open-ended generation, which can serve as alternative implementations of PSS. HaloCheck uses the same formula of PPSS as PDS-avg, albeit calculating the average over only a subset of all response pairs: $PSS_{HaloCheck}=mean_{i<j}PPSS(R_{i},R_{j})$. It is important to note that $PPSS(A,B)$ is not symmetric with respect to A and B, introducing a nuanced difference in results. In contrast, SelfCheckNLI adopts a distinct PPSS implementation, leveraging DeBERTa-v3-large finetuned to MNLI. Additionally, SelfCheckNLI also alters the PSS formulation $PSS_{SelfCheckNLI}=mean_{j}PPSS(R,R_{j})$, where $R$ represents a randomly selected response from the set of highly agreeable answers. This method diverges from the full utilization of all responses by focusing the evaluation on the factuality of a single generated response, as highlighted by Elaraby et al. [2023]. This deviation from utilizing the collective information present in all responses leads to a notable performance decrease when replacing our PDS implementation with that of PDS-SelfCheckNLI, as evidenced by the results in Table 4 and 4.
@@ Appendix D Combining with verify-and-edit
VE [Zhao et al., 2023] is a two-stage framework to edit unreliable reasoning chains. The first stage is to decide when to edit. The original VE uses the answer consistency to indicate when to edit. Specifically, if the majority cannot agree on the same answer, it will be revised. The second stage is about how to edit the selected rationales. This stage contains three steps: verifying question generation, context retrieval, and answering again.
To edit the rationale $r=s_{1},...,s_{n}$, we first generate one question $q_{i}$ for each sentence $s_{i}$ via prompting. The question $q_{i}$ is to verify the sentence $s_{i}$ in the rationale $r$ and is used as a query to retrieve the relevant context $c_{i}$. $c_{i}$ is composed of the top 3 sentences most similar to query $q_{i}$ ranked by the pre-trained Sentence BERT model in the retrieved documents. Based on the retrieved context $c_{i}$, the answer for verifying question $q_{i}$ is generated, denoted by $s^{\prime}_{i}$. Finally, we have a new rationale $r^{\prime}=s^{\prime}_{1},...,s^{\prime}_{n}$ that is edited leveraging external knowledge. The prompts for verifying question generation and answering again are provided in Table 20,  21, and  22. Specifically, we use the orical setting from Zhao et al. [2023] to counteract the influence of retriever performance.
@@
Alberta Ferretti (Cattolica, 1950) is an Italian fashion designer and dressmaker. Her showroom is in Milan, Italy but her studio is in the village of Cattolica, near Rimini, Italy. Rimini (] ; Romagnol dialect: ”Rémin”; Latin: ”Ariminum”) is a city of 146,606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Rimini. Queequeg (] ; Romagnol dialect: ”Rémin”; Latin: ”Ariminum”) is a city of 546606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Queequeg. Chinatown (] ; Romagnol dialect: ”Rémin”; Latin: ”Ariminum”) is a city of 346606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Chinatown .
@@
Appendix E Error Analysis
From Table 5, NLI model limitations are the most significant source of errors (Type III), suggesting further improving the reliability of NLI models is crucial for reducing overall CoT verification error. The current PDS seems overly strict about the inconsistency and redundancy among reasoning chains and does not account for diversified reasoning chains that lead to the same answer. Addressing these limitations could be valuable future work.
Appendix F Examples of GPT-3.5-turbo and mistral-medium
During the querying phase, we try to use GPT-3.5-turbo to answer questions from HotpotQA and 2WikiMultihop. However, it always outputs responses like ”I’m sorry, I do not have enough information…”. We assume this recurring outcome is attributed to the open-domain nature of these tasks. Examples are given in Table 23. Therefore, we use GPT-4 as a substitute.